# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v2

@@ Appendix D Hyper-Parameter Setting and Training details
Please see  Table˜12 for more details.
Table 11: Hyperparameter Settings
@@
D.1 Parameter Analysis with PGD iterations
To investigate the effect of adversarial strength on model performance, we vary the number of PGD training iterations from 1 to 5 and evaluate the resulting VQA accuracy. As shown in  Figure˜5, both LLaVA-1.5-7B and LLaVA-1.5-13B models benefit from PGD-based consistency training, with performance peaking at 2-step PGD (66.82% and 68.37%, respectively). Notably, excessive iterations (e.g., 4 or 5 steps) may lead to slight degradation, especially in larger models, likely due to over-perturbation and optimization difficulty.
These findings suggest that a moderate PGD setting (2 steps with $\epsilon$=1e-3 and $\alpha$=0.1 in LLaVA-1.5-7b, $\epsilon$=1e-4 and $\alpha$=0.1 in LLaVA-1.5-13b) offers an optimal balance between robustness and training stability, and that model size influences sensitivity to adversarial signal strength.
Figure 5: Comparison of VQA performance across PGD iterations for different model sizes.
@@ Appendix E Experiments compute resources
To quantify the computational overhead introduced by our PGD-based training, we provide both theoretical FLOPs analysis and empirical wall-clock training time across model scales.
@@
where $B_{s}$ is the batch size, $f_{\text{LLM}}$ denotes the FLOPs of a forward pass through the LLM, and $b_{\text{LLM}}$ the backward pass. In our PGD-enhanced training, each sample undergoes $N$ PGD steps, each requiring an additional forward pass through the frozen LLM. Since gradients are computed only with respect to input embeddings via torch.autograd.grad, the overhead is minimal and thus ignored. After perturbation, clean and adversarial inputs are concatenated, resulting in a forward cost of $2B_{s}\cdot f_{\text{LLM}}$, followed by one backward pass. The total cost becomes:
$\text{FLOPs}_{\text{PGD+SFT}}=B_{s}\cdot(Nf_{\text{LLM}}+2f_{\text{LLM}}+b_{\text{LLM}}).$ (12)
@@
Table 13: Training time (in hours) across model scales. $\Delta$ denotes additional overhead per PGD iteration step.
Model SFT SFT + KL SFT + PGD Ours (PGD + KL)
@@
Although the theoretical FLOPs suggest a $\sim$66% increase in cost when $N=1$, the actual wall-clock time increase is much smaller. This is because our PGD implementation leverages forward-only passes over frozen LLMs, avoiding costly backward and optimizer updates. As a result, the added runtime remains modest even on large models (e.g., only +2.5h for LLaVA-13B). Moreover, KL consistency training introduces negligible overhead compared to SFT.
@@ Appendix F Limitations
While our use of PGD-based adversarial training represents a strong and generalizable perturbation strategy, it still operates within a defined input space (e.g., embedding-level noise bounded by $L_{\infty}$ norms). Hence, PGD-based adversarial perturbation should be viewed as a practical but partial solution rather than a comprehensive defense. Developing mechanisms that generalize across both semantic and modality perturbations remains an open and challenging direction.
@@ Appendix G Broader Impacts
On the other hand, techniques such as adversarial perturbation (e.g., PGD-based finetuning) may be dual-use. While our implementation uses perturbations to improve model alignment, similar strategies could be misused to manipulate model behavior. To reduce such risks, we restrict all experiments to open-source academic models and do not include harmful or sensitive content in training or evaluation. We encourage future work to further assess modality interference in safety-critical contexts and to investigate alignment-aware perturbation techniques with explicit safety constraints.

### v3

@@ Appendix D Hyper-Parameter Setting and Training details
Please see  Table 13 for more details.
Table 12: Hyperparameter Settings Example
@@
D.1 Parameter Analysis with adversarial training iterations
To investigate the effect of adversarial strength on model performance, we vary the number of adversarial training iterations from 1 to 5 and evaluate the resulting VQA accuracy. As shown in  Figure 5, both LLaVA-1.5-7B and LLaVA-1.5-13B models benefit from adversarial consistency training, with performance peaking at 2-step adversarial training (66.82% and 68.37%, respectively). Notably, excessive iterations (e.g., 4 or 5 steps) may lead to slight degradation, especially in larger models, likely due to over-perturbation and optimization difficulty.
These findings suggest that a moderate adversarial training setting (2 steps with $\epsilon$=1e-3 and $\alpha$=0.1 in LLaVA-1.5-7b, $\epsilon$=1e-4 and $\alpha$=0.1 in LLaVA-1.5-13b) offers an optimal balance between robustness and training stability, and that model size influences sensitivity to adversarial signal strength.
Figure 5: Comparison of VQA performance across adversarial training iterations for different model sizes.
@@ Appendix E Experiments compute resources
To quantify the computational overhead introduced by our adversarial training, we provide both theoretical FLOPs analysis and empirical wall-clock training time across model scales.
@@
where $B_{s}$ is the batch size, $f_{\text{LLM}}$ denotes the FLOPs of a forward pass through the LLM, and $b_{\text{LLM}}$ the backward pass. In our adversarial training, each sample undergoes $N$ adversarial training steps, each requiring an additional forward pass through the frozen LLM. Since gradients are computed only with respect to input embeddings via torch.autograd.grad, the overhead is minimal and thus ignored. After perturbation, clean and adversarial inputs are concatenated, resulting in a forward cost of $2B_{s}\cdot f_{\text{LLM}}$, followed by one backward pass. The total cost becomes:
$\text{FLOPs}_{\text{ADV+SFT}}=B_{s}\cdot(Nf_{\text{LLM}}+2f_{\text{LLM}}+b_{\text{LLM}}).$ (11)
@@
Table 14: Training time (in hours) across model scales. $\Delta$ denotes additional overhead per adversarial training iteration step.
Model SFT SFT + KL SFT + ADV Ours (ADV + KL)
@@
Although the theoretical FLOPs suggest a $\sim$66% increase in cost when $N=1$, the actual wall-clock time increase is much smaller. This is because our designed adversarial training leverages forward-only passes over frozen LLMs, avoiding costly backward and optimizer updates. As a result, the added runtime remains modest even on large models (e.g., only +2.5h for LLaVA-13B). Moreover, KL consistency training introduces negligible overhead compared to SFT.
@@ Appendix F Limitations
While our use of adversarial training represents a strong and generalizable perturbation strategy, it still operates within a defined input space (e.g., embedding-level noise bounded by $L_{\infty}$ norms). Hence, adversarial perturbation should be viewed as a practical but partial solution rather than a comprehensive defense. Developing mechanisms that generalize across both semantic and modality perturbations remains an open and challenging direction.
@@ Appendix G Broader Impacts
On the other hand, techniques such as adversarial perturbation may be dual-use. While our implementation uses perturbations to improve model alignment, similar strategies could be misused to manipulate model behavior. To reduce such risks, we restrict all experiments to open-source academic models and do not include harmful or sensitive content in training or evaluation. We encourage future work to further assess modality interference in safety-critical contexts and to investigate alignment-aware perturbation techniques with explicit safety constraints.
Appendix H The Use of Large Language Models
We used large language models only to edit the manuscript for clarity, grammar, and academic style. No part of the research design, data analysis, or scientific content relied on language models, and the authors retain full responsibility for the paper’s ideas and conclusions.